A woman has $5,000 invested at 8% annual interest. At what rate must she invest an additional $10,000 so that her annual income from both investments is 9% of her total investment?
Answer by TimothyLamb(4379)

x = rate at which 10000 is invested
---
5000*0.08 + 10000x = 15000*0.09
10000x = 15000*0.09 - 5000*0.08
x = (15000*0.09 - 5000*0.08)/10000
x = 0.095
---
answer:
invest the additional %10,000 at 9.5%
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---